Please hep me solve this 2 questions:
1.If f(x)= x^2, when x>0 and c>1, then f(x^c)= ?
2. 2log10x-log10y=?
Thanks

Answer by stanbon(75887)

1.If f(x)= x^2, when x>0 and c>1, then f(x^c)= ?
---
f(x^c) = (x^c)^2 = x^(2c)
=================================
2. 2log10x-log10y=?
---
= log10(x^2)-log10(y)
-----
= log10[x^2/y]
=====================
Cheers,
Stan H.